# A lack of LAEs within 5Mpc of a luminous quasar in an overdensity at z=6.9: potential evidence of quasar negative feedback at protocluster scales Key Words.: Cosmology: reionization – Galaxies: high-redshift – Galaxies: quasars: individual:VIK J2348–30541 Introduction

### v1

@@
High-redshift quasars are thought to live in the densest regions of space which should be made evident by an overdensity of galaxies around them. However, campaigns to identify these overdensities through the search of Lyman Break Galaxies (LBGs) and Lyman emitters (LAEs) have had mixed results. These may be explained by either the small field of view of some of the experiments, the broad redshift ranges targeted by LBG searches, and by the inherent large uncertainty of quasar redshifts estimated from UV emission lines, which makes it difficult to place the Ly- emission line within a narrowband filter. Here we present a three square degree search ( pMpc) for LAEs around the quasar VIK J2348–3054 using the Dark Energy CAMera (DECam), housed on the 4m Blanco telescope, finding 38 LAEs. The systemic redshift of VIK J2348–3054 is known from ALMA [CII] observations and place the Ly- emission line of companions within the NB964 narrowband of DECam. This is the largest field of view LAE search around a quasar conducted to date. We find that this field is 10 times more overdense when compared to the Chandra Deep-Field South, observed previously with the same instrumental setup as well as several combined blank fields. This is strong evidence that VIK J2348–3054 resides in an overdensity of LAEs over several Mpc. Surprisingly, we find a lack of LAEs within 5 physical Mpc of the quasar and take this to most likely be evidence of the quasar suppressing star formation in its immediate vicinity. This result highlights the importance of performing overdensity searches over large areas to properly assess the density of those regions of the Universe.
@@
Another observational effect, which is often not considered when searching for overdensities around quasars, is the method used to determine the quasar redshift. Most studies adopted quasar redshifts that have been measured using rest-frame UV lines. However, with the advent of ALMA, it is now clear that the redshifts from UV can be offset by thousands of km s from the host galaxy’s systemic redshifts (e.g., Decarli et al., 2018; Schindler et al., 2020; Díaz-Santos et al., 2021). This offset can often be large enough to shift the Ly- emission outside of a narrowband filter in LAE searches, potentially resulting in false non-detections of overdensities. To date, only Ota et al. (2018), who searched over a 0.2 square degree area, has performed a LAE search using a [CII] redshift instead of rest-frame UV. While they did not find an overdensity, it is worth noting that the Ly- line in their case, was near the edge of their narrowband filter.
@@
The size of the search area may also affect the identification of overdensities. Protocluster areas are expected to be on scales of (Overzier et al., 2009; Balmaverde et al., 2017). Chiang et al. (2017) used a semi-analytical model anchored on the Millennium simulation and found that, on average, haloes with radii cMpc at z=0 are extended over 10 cMpc ( pMpc) at z=7; Overzier et al. (2009) suggests that protoclusters typically extend over 25 cMpc but can reach up to 75 cMpc. However, most studies probe areas significantly smaller than this (e.g., Bañados et al., 2013; Mazzucchelli et al., 2017a).
@@
Accounting for these observational biases would require 1) a quasar with a known, systematic redshift, 2) a telescope and instrument with a large FoV, and 3) a narrowband filter at the correct wavelength range to capture Lyman alpha at the systemic velocity. By coincidence, this is exactly the case for the quasar VIK 2348–3054 (RA = 23:48:33.34, Dec = 30:54:10.0; Venemans et al., 2013), which has a confirmed [CII] redshift of (Venemans et al., 2016), and whose Ly- emission falls within the response of the narrowband filter NB964 (Zheng et al., 2019) available for the Dark Energy Camera (DECam, Flaugher et al., 2015) on the 4m Blanco telescope at Cerro Tololo Inter-American Observatory (CTIO), which has a 3 square degree FoV. The narrowband filter NB964 was designed for the Lyman Alpha Galaxies in the Epoch of Reionization (LAGER, Zheng et al., 2017) survey.
@@
In this paper, we take full advantage of this serendipitous combination of telescope, instrument, and target and perform a LAE search around the quasar VIK J2348–3054, with a bolometric luminosity and black hole mass of erg s and respectively (see Mazzucchelli et al., 2017b). The remainder of the paper is organized as follows: Section 2 describes the observational setup adopted to perform the search. We present the results in Section 3, and in Section 4 we discuss the implications of these results and offer an explanation as to the current tension within the literature. We present our conclusions in Section 5.
@@
Throughout this paper we adopt a vanilla, flat CDM cosmology with , and km s Mpc. At the Universe is 0.753 Gyr old and the transverse scales are 24 arcseconds/cMpc and 189 arcseconds/pMpc. All magnitudes are in the AB system.
@@ 2.1 DECam Observations of VIK J2348–3054
We observed the area around VIK J2348–3054 with DECam housed on the 4m Blanco telescope at the CTIO (PROPID: 2021B-0905). The observing run was split into two parts: the first took place over four half-nights from the UT 2021-08-30 to the UT 2021-09-03, while the second occurred over two half-nights on UT 2021-10-20 and 2021-10-21. From the combined seven half-nights of observations we obtained integrations of 5.8, 8.6, and 14.2 hours in the i-band ( Å; = 1470 Å), z-band ( Å; Å), and NB964 narrowband ( Å; = 94 Å), respectively. The filter transmission curves are shown in Figure 1.
This particular choice of filters allow us to search for LAEs within a redshift range of , making it ideal for a search around VIK J2348–3054 at (Venemans et al., 2016). DECam has 64 individual CCDs with numerous gaps between them. To account for this and obtain the best possible data quality, we implemented a dithering pattern during the observations, randomly shifting the pointing position for each set of exposures.
Each band was stacked using the DECam community pipeline (Valdes et al., 2014), then astronomically aligned to the GAIA DR2 catalog (Gaia Collaboration et al., 2018) and finally trimmed to cover the same area in all bands.
We masked several obvious CCD artifacts, which included stellar haloes and image “ringing”, i.e., clusters of two or three unrealistically bright and/or negative pixels, by eye on a case by case basis. We also masked noisy edges by creating a polygon region within the borders of the combined images. The final usable field of view (FoV) is 2.87 deg. At the redshift of VIK J2348–3054 this area is equivalent to 1034 pMpc
@@ 2.1.1 Photometric Calibration
Previous DECam NB964 filter calibrations have often used Spectral Energy Distribution (SED) fitting of bright stars (either A, B or F0 stars) identified in the field (Wold et al., 2019a; Hu et al., 2019; Khostovan et al., 2020). However, Wold et al. (2019a) showed that calculating zero points assuming a linear relationship between the and colors, for bright, point like sources, is equivalent to SED fitting. We show this relationship for our data in Figure 2 by plotting the color-color diagram of sources which were identified as point sources by SExtractor (Bertin & Arnouts, 1996), and had z-band, y-band and i-band colors in the Panoramic Survey Telescope and Rapid Response System (Pan-STARRS1) first data release (Chambers et al., 2016). We therefore photometrically calibrated the NB964 filter using the method by Wold et al. (2019b, 2022), assuming a linear relation between the PS1 and DECam magnitudes for point sources:
where and are the and Pan-STARRS1 magnitudes respectively, NB964 is the instrumental narrowband magnitude in DECam, and and ZPT are the slope and zero point which are needed to transform the magnitudes. The and ZPT values are solved by fitting a best-fit straight line, shown in Figure 2, and where determined to be and respectively. The shaded area represents the value of the linear fit.
The photometric transformations for the DECam broadband filters are available via the Dark Energy Survey data management (Abbott et al., 2021) who used (amongst other surveys) the Pan-STARRS1 survey 111https://des.ncsa.illinois.edu/releases/dr2/dr2-docs/dr2-transformations. Following their work, we photometrically calibrated the broadband filters using the following transformations:
@@ 2.1.2 Depth
The limiting magnitudes were determined by fitting an exponential function to the relationship between detected sources’ magnitudes and their signal-to-noise ratios (SNRs), given by:
@@ 2.2 DECam Observations of the CDFS LAGER field
We use the Lyman Alpha Galaxies in the Epoch of Reionization (LAGER) CDFS field observations from Hu et al. (2019), for comparison since it does not contain a known quasar or any other feature that would make it unusual at the redshift of VIK J2348–3054. This field was observed with the same telescope, instrument, and filters that we used to search for LAEs around VIK J2348–3054. The CDFS images have several CCD artifacts as well as large areas of low SNR, mainly within the chip gaps due to insufficient dithering. We manually masked CCD artifacts and automatically masked the chip gap regions in each image using the corresponding weights image. The effective area available for source detection after masking was deg, or pMpc at the redshift of VIK J2348–3054.
@@
While the narrowband depths are similar between the observations around VIK J2348–3054 and CDFS, the broadbands in the latter are about 2 magnitudes deeper (see Hu et al., 2019). To create a comparison sample with the same photometric properties, it is necessary to degrade the CDFS photometry to mimic our depths. To do this we create a source catalog using SExtractor in the same way that we did for our data and then assign new errors to every detected source based on the fitted exponential functions from the depth calculations for our data (discussed §2.1.3) resulting in a degraded magnitude error, . To replicate the expected scatter due to a lower depth image, we randomly assign a new source magnitude which is drawn from a normal distribution with a mean equal to the magnitude determined in the CDFS image and a standard deviation equal to , where is the original magnitude error determined by SExtractor in the CDFS image.
@@ 2.3 Selection Criteria
A typical LAE has very strong Ly- emission and a faint continuum (Malhotra & Rhoads, 2002; Ouchi et al., 2020). We plot a synthetic LAE spectrum at the redshift of VIK J2348–3054, along with the filter curves in Figure 1. The standard IGM extinction from Madau (1995) was assumed. The line was made by assuming a Gaussian profile and adopting a 200 km s FWHM value and an equivalent width of 50 Å. The luminosity is the L and was derived using the luminosity function, reported in Matthee et al. (2015). As expected, the strong Ly- emission is contained within the NB964 filter, with practically no flux in the i-band, while the z-band is dominated by the source continuum. Given this, we adopt a similar selection criteria to that used by Bañados et al. (2013) and Mazzucchelli et al. (2017a), namely:
@@
To further justify our selection we take the synthetic LAE spectrum from Figure 1, and derive the expected colors using the python package synphot222Version 1.2.1 in the redshift range . This color-color track is shown in Figure 4 for equivalent widths of 10Å and 50Å. The Figure shows that the LAEs meet the criteria at the targeted redshift for all EWs of Lyman alpha considered in Figure 3. At the same time we tested two types of interlopers using the spectral energy distributions provided in Polletta et al. (2007): M82, and QSO2. These represent a star forming galaxy and a type-2 quasar respectively. We plot their color-color tracks from in green and blue and include no dust extinction and E(B-V) = 1. Neither of these two low-z interlopers enter our color selection (indicated as the top right green box in Figure 4). This justifies our color selections, demonstrating that these choices are optimal to select Ly- emitters around the quasar redshift.
@@ 3.1 LAE Candidates
We find 39 sources sources in the VIK J2348–3054 field which meet our selection criteria: 38 LAE candidates as well as the quasar itself. The 20 pix 20 pix (10.8” 10.8”) postage stamp cut outs of all the candidates are presented in Figure 6. We also indicate the magnitudes in the respective bands as well as their SNRs. While all candidates formally meet our selection criteria (see §2.3), we do note that three candidates (LAE-25, LAE-30, and LAE-38) seem to have obvious i-band detections despite SExtractor assigning them very low SNR values. By contrast, only two candidates were identified in the CDFS images after the photometry degradation (see §2.3).
Our candidate properties are presented in Table 1, including their positions, separation from VIK J2348–3054, i-band, z-band, and NB964 magnitudes, Ly- luminosities, and star formation rates (see sections 3.2 and 3.3).

### v2

@@
High-redshift quasars are thought to live in the densest regions of space which should be made evident by an overdensity of galaxies around them. However, campaigns to identify these overdensities through the search of Lyman Break Galaxies (LBGs) and Lyman emitters (LAEs) have had mixed results. These may be explained by either the small field of view of some of the experiments, the broad redshift ranges targeted by LBG searches, and by the inherent large uncertainty of quasar redshifts estimated from UV emission lines, which makes it difficult to place the Ly- emission line within a narrowband filter. Here we present a three square degree search ( pMpc2) for LAEs around the quasar VIK J2348–3054 using the Dark Energy CAMera (DECam), housed on the 4m Blanco telescope, finding 38 LAEs. The systemic redshift of VIK J2348–3054 is known from ALMA [CII] observations and place the Ly- emission line of companions within the NB964 narrowband of DECam. This is the largest field of view LAE search around a quasar conducted to date. We find that this field is 10 times more overdense when compared to the Chandra Deep-Field South, observed previously with the same instrumental setup as well as several combined blank fields. This is strong evidence that VIK J2348–3054 resides in an overdensity of LAEs over several Mpc. Surprisingly, we find a lack of LAEs within 5 physical Mpc of the quasar and take this to most likely be evidence of the quasar suppressing star formation in its immediate vicinity. This result highlights the importance of performing overdensity searches over large areas to properly assess the density of those regions of the Universe.
@@
Another observational effect, which is often not considered when searching for overdensities around quasars, is the method used to determine the quasar redshift. Most studies adopted quasar redshifts that have been measured using rest-frame UV lines. However, with the advent of ALMA, it is now clear that the redshifts from UV can be offset by thousands of km s-1 from the host galaxy’s systemic redshifts (e.g., Decarli et al., 2018; Schindler et al., 2020; Díaz-Santos et al., 2021). This offset can often be large enough to shift the Ly- emission outside of a narrowband filter in LAE searches, potentially resulting in false non-detections of overdensities. To date, only Ota et al. (2018), who searched over a 0.2 square degree area, has performed a LAE search using a [CII] redshift instead of rest-frame UV. While they did not find an overdensity, it is worth noting that the Ly- line in their case, was near the edge of their narrowband filter.
@@
The size of the search area may also affect the identification of overdensities. Protocluster areas are expected to be on scales of (Overzier et al., 2009; Balmaverde et al., 2017). Chiang et al. (2017) used a semi-analytical model anchored on the Millennium simulation and found that, on average, haloes with radii cMpc at z=0 are extended over 10 cMpc ( pMpc) at z=7; Overzier et al. (2009) suggests that protoclusters typically extend over 25 cMpc but can reach up to 75 cMpc. However, most studies probe areas significantly smaller than this (e.g., Bañados et al., 2013; Mazzucchelli et al., 2017a), although there has been an effort made in the submm regime, specifically by Li et al. (2023) to search for submm galaxies around these quasars over larger areas.
@@
Accounting for these observational biases would require 1) a quasar with a known, systemic redshift, 2) a telescope and instrument with a large FoV, and 3) a narrowband filter at the correct wavelength range to capture Lyman alpha at the systemic velocity. By coincidence, this is exactly the case for the quasar VIK 2348–3054 (RA = 23:48:33.34, Dec = 30:54:10.0; Venemans et al., 2013), which has a confirmed [CII] redshift of (Venemans et al., 2016), and whose Ly- emission falls within the response of the narrowband filter NB964 (Zheng et al., 2019) available for the Dark Energy Camera (DECam, Flaugher et al., 2015) on the 4m Blanco telescope at Cerro Tololo Inter-American Observatory (CTIO), which has a 3 square degree FoV. The narrowband filter NB964 was designed for the Lyman Alpha Galaxies in the Epoch of Reionization (LAGER, Zheng et al., 2017) survey.
@@
In this paper, we take full advantage of this serendipitous combination of telescope, instrument, and target and perform a LAE search around the quasar VIK J2348–3054, with a bolometric luminosity and black hole mass of erg s-1 and respectively, both of which are very close to the median values of all known quasars at this redshift (see Table 8 in Mazzucchelli et al. (2017b) and Table 1 in Mazzucchelli et al. (2023). The remainder of the paper is organized as follows: Section 2 describes the observational setup adopted to perform the search. We present the results in Section 3, and in Section 4 we discuss the implications of these results and offer an explanation as to the current tension within the literature. We present our conclusions in Section 5.
@@
Throughout this paper we adopt a vanilla, flat CDM cosmology with , and km s-1 Mpc-1. At the Universe is 0.753 Gyr old and the transverse scales are 24 arcseconds/cMpc and 189 arcseconds/pMpc. All magnitudes are in the AB system.
@@ 2.1 DECam Observations of VIK J2348–3054
We observed the area around VIK J2348–3054 with DECam housed on the 4m Blanco telescope at the CTIO (PROPID: 2021B-0905). The observing run was split into two parts: the first took place over four half-nights from the UT 2021-08-30 to the UT 2021-09-03 where the moon did not rise during the observations, while the second occurred over two half-nights on UT 2021-10-20 and 2021-10-21 during full-moon. All observations were done under acceptable observing conditions and all frames were used. From the combined seven half-nights of observations we obtained integrations of 5.8, 8.6, and 14.2 hours in the i-band ( Å; = 1470 Å), z-band ( Å; Å), and NB964 narrowband ( Å; = 94 Å), respectively. The filter transmission curves are shown in Figure 1.
This particular choice of filters allow us to search for LAEs within a redshift range of 6.89 ¡ z ¡ 6.97, making it well suited for a search around VIK J2348–3054 at z = 6.9018 (Venemans et al., 2016). Given that the Ly- line tends to be redshifted from the galaxy’s systemic velocity by 200 km s-1 (Verhamme et al., 2018), we can detect galaxies with peculiar velocities ranging from -700 km s-1 to 2250 km s-1 from the quasar’s systemic velocity. Sources around high-z quasars and in protocluster environments can have peculiar velocities within 1000 km s-1 (Overzier et al., 2009). This means that we should detect all sources associated with the protocluster environment with redshifted peculiar velocities, but may miss some sources with extreme blue shifted peculiar velocities.
DECam has 64 individual CCDs with numerous gaps between them. To account for this and obtain the best possible data quality, we implemented a dithering pattern during the observations, randomly shifting the pointing position within a 120 arcsecond box for each set of exposures. The pointing was also offset globally by 5 arcminutes in order to ensure that the quasar was not centered on a chip gap111The full observation script-generator is available at https://github.com/TrystanScottLambert/DECam_Photometry/blob/main/scriptmaker.py.
Each band was stacked using the DECam community pipeline (Valdes et al., 2014). The DECam Community Pipeline provides the following operations: linearity correction, bias subtraction, dome, photometric, and dark sky flat fielding, bad pixel masking, masking of saturation effects and cosmic rays, an astrometric solution, removal of background patterns including stray light, pupil reflection, and fringing, and projection to a standard tangent plane. For multiple, dithered exposures a co-add is produced which includes transient masking. Three methods for masking cosmic rays and other transients were used in this work: the “maskcosmics” algorithm, from the Dark Energy Survey, using the traditional comparison to the PSF; an algorithm that identifies “long” cosmic rays based on the ratio of boundary to interior pixels from the IRAF segmentation and cataloging package; individual exposures are compared to a median by image differencing detection; and finally, statistical outlier rejection may be used when making the final stack.
After being stacked the images were astronomically aligned to the Gaia DR2 catalog (Gaia Collaboration et al., 2018) and finally trimmed to cover the same area in all bands.
We masked several obvious CCD artifacts, which included stellar haloes and image “ringing”, i.e., clusters of two or three unrealistically bright and/or negative pixels, by eye on a case by case basis. We also masked noisy edges by creating a polygon region within the borders of the combined images. The final usable field of view (FoV) is 2.87 deg2. At the redshift of VIK J2348–3054 this area is equivalent to 1034 pMpc2
@@ 2.1.1 Photometric Calibration
Previous DECam NB964 filter calibrations have often used Spectral Energy Distribution (SED) fitting of bright stars (either A, B or F0 stars) identified in the field (Wold et al., 2019; Hu et al., 2019; Khostovan et al., 2020). However, Wold et al. (2019) showed that calculating zero points assuming a linear relationship between the and colors, for bright, point like sources, is equivalent to SED fitting. We show this relationship for our data in Figure 2 by plotting the color-color diagram of sources which were identified as point sources by SExtractor (Bertin & Arnouts, 1996), and had z-band, y-band and i-band colors in the Panoramic Survey Telescope and Rapid Response System (Pan-STARRS1) first data release (Chambers et al., 2016). We therefore photometrically calibrated the NB964 filter using the method by Wold et al. (2019, 2022), assuming a linear relation between the PS1 and DECam magnitudes for point sources:
where and are the and Pan-STARRS1 magnitudes respectively, NB964inst is the instrumental narrowband magnitude in DECam, and and ZPT are the slope and zero point which are needed to transform the magnitudes. The and ZPT values are solved by fitting a best-fit straight line to 685 identified point sources, shown in Figure 2. They where determined to be and respectively. The shaded area represents the value of the linear fit.
The photometric transformations for the DECam broadband filters are available via the Dark Energy Survey data management (Abbott et al., 2021) who used (amongst other surveys) the Pan-STARRS1 survey 222https://des.ncsa.illinois.edu/releases/dr2/dr2-docs/dr2-transformations. Following their work, we photometrically calibrated the broadband filters using the following transformations:
@@ 2.1.2 Depth
The limiting magnitudes were determined by fitting an exponential function to the relationship between detected sources’ magnitudes and their signal-to-noise ratios (S/N), given by:
@@ 2.2 DECam Observations of the CDFS LAGER field
We use the Lyman Alpha Galaxies in the Epoch of Reionization (LAGER) CDFS field observations from Hu et al. (2019), for comparison since it does not contain a known quasar or any other feature that would make it unusual at the redshift of VIK J2348–3054. This field was observed with the same telescope, instrument, and filters that we used to search for LAEs around VIK J2348–3054. The CDFS images have several CCD artifacts as well as large areas of low S/N, mainly within the chip gaps due to insufficient dithering. We manually masked CCD artifacts and automatically masked the chip gap regions in each image using the corresponding weights image. The effective area available for source detection after masking was deg2, or pMpc2 at the redshift of VIK J2348–3054.
@@
While the narrowband depths are similar between the observations around VIK J2348–3054 and CDFS, the broadbands in the latter are about 2 magnitudes deeper (see Hu et al., 2019). To create a comparison sample with the same photometric properties, it is necessary to degrade the CDFS photometry to mimic our depths. To do this we create a source catalog using SExtractor in the same way that we did for our data and then assign new errors to every detected source based on the fitted exponential functions from the depth calculations for our data (discussed Sect. 2.1.3) resulting in a degraded magnitude error, . To replicate the expected scatter due to a lower depth image, we randomly assign a new source magnitude which is drawn from a normal distribution with a mean equal to the magnitude determined in the CDFS image and a standard deviation equal to , where is the original magnitude error determined by SExtractor in the CDFS image.
@@ 2.3 Selection Criteria
A typical LAE has very strong Ly- emission and a faint continuum (Malhotra & Rhoads, 2002; Ouchi et al., 2020). We plot a synthetic LAE spectrum at the redshift of VIK J2348–3054, along with the filter curves in Figure 1. The standard IGM extinction from Madau (1995) was assumed. The line was made by assuming a Gaussian profile and adopting a 200 km s-1 FWHM value and an equivalent width of 50 Å. The luminosity is the L∗ and was derived using the luminosity function, reported in Matthee et al. (2015). As expected, the strong Ly- emission is contained within the NB964 filter, with practically no flux in the i-band, while the z-band is dominated by the source continuum. Given this, we adopt a similar selection criteria to that used by Bañados et al. (2013) and Mazzucchelli et al. (2017a), namely:
@@
To further justify our selection we take the synthetic LAE spectrum from Figure 1, and derive the expected colors using the python package synphot333Version 1.2.1 in the redshift range . This color-color track is shown in Figure 4 for equivalent widths of 10Å and 50Å. The Figure shows that the LAEs meet the criteria at the targeted redshift for all EWs of Lyman alpha considered in Figure 3. At the same time we tested two types of interlopers using the spectral energy distributions provided in Polletta et al. (2007): M82, and QSO2. These represent a star forming galaxy and a type-2 quasar respectively. We plot their color-color tracks from in green and blue and include no dust extinction and E(B-V) = 1. Neither of these two low-z interlopers enter our color selection (indicated as the top right green box in Figure 4). This justifies our color selections, demonstrating that these choices are optimal to select Ly- emitters around the quasar redshift.
@@ 3.1 LAE Candidates
We find 39 sources sources in the VIK J2348–3054 field which meet our selection criteria: 38 LAE candidates as well as the quasar itself. The 20 pix 20 pix (10.8” 10.8”) postage stamp cut outs of all the candidates are presented in Figure 11 in the appendix. We also indicate the magnitudes in the respective bands as well as their S/N. While all candidates formally meet our selection criteria (see Sect. 2.3), we do note that three candidates (LAE-25, LAE-30, and LAE-38) seem to have obvious i-band detections despite SExtractor assigning them very low S/N values. By contrast, only two candidates were identified in the CDFS images after the photometry degradation (see Sect. 2.3).
Our candidate properties are presented in Table 2 in the appendix, including their positions, separation from VIK J2348–3054, i-band, z-band, and NB964 magnitudes, Ly- luminosities, and star formation rates (see sections 3.2 and 3.3). VIK2348–3054 also met the selection criteria and is therefore also presented in Table 1. The z magnitude determined for the quasar is nearly 0.8 mag fainter than the one reported in Venemans et al. (2013), upon discovery, who measured a Gunn z filter magnitude of on the 3.8m New Technology Telescope (NTT). The slight difference between the two is easily explainable by the difference in filter sets between DECam and NTT; the z-band in DECam is much redder than in EFOCS2 (because of the lower efficiency of the CCD at long wavelengths). This is also notable in Table 3 of Mazzucchelli et al. (2017b) where the PS1 z-band magnitudes differ from the DECam z-band photometry by substantial amounts due to the filter curve. Additionally, our DECam z value is also expected given the relationship between previous quasars’ DECam z magnitudes and their J band magnitudes (see Table in 3 in Mazzucchelli et al., 2017a) which have comparable redshifts and luminosities to VIKJ2348-3054.